# Deep learning for predicting the occurrence of tipping points

### v1

@@ Introduction
A classical mathematical tool to understand tipping points is bifurcation theory which focuses on how dynamical systems undergo sudden qualitative changes as a parameter crosses a threshold(7). There are basically two classes of methods based on bifurcation theory to deal with tipping points prediction. The first class of methods is using lag-1 autocorrelation which is based on “critical slowing down”, a paramount clue of whether a tipping point is approached(5, 8). Such methods are widely used in various complex systems, including degenerate fingerprinting(9), BB method(10) and ROSA(11). The second class of methods is using dynamical eigenvalue (DEV)(12) which is based on Takens embedding theorem(13). Yet, the success of the existing methods based on bifurcation theory are impaired by two fundamental limitations. First, these methods are not applicable for irregularly-sampled time series data which are commonly observed from various scientific fields, such as geoscientific measurements(14, 15), medical observations(16) and biological systems(17). Second, the performance of existing methods based on bifurcation theory in prediction accuracy is affected by several approximations used in these methods. One such approximation is that only the first-order term of the dynamical systems is considered in these methods while the impact of the higher-order terms is ignored. However, the higher-order terms become significant for prediction when a tipping point is approached(18) (see Supplementary Information section 1 for details). Besides, in the approximation of fast-slow systems(19), the delay on the bifurcation-tipping due to the changing rate of the bifurcation parameter is ignored(20, 21) (see Supplementary Information section 2.1 and 2.2 for details).
@@
Recent studies show that some machine learning based techniques have been used as effective early warning signals of tipping points by learning generic features of bifurcation(18, 22, 23, 24). However, they can not be used for predicting where the tipping points occur. A machine learning framework based on reservoir computing has been developed for tipping points prediction(25, 26, 27). But this algorithm is only effective for nonlinear dynamical systems with chaotic attractors and requires a large amount of data from phase space for training.
@@
To circumvent above limitations of existing methods for tipping points prediction, we develop a deep learning (DL) algorithm based on 2D CNN-LSTM architecture. This DL algorithm only requires the time series of the state and that of the bifurcation parameter from a single variable of a system. The 2D CNN layer is used to extract local features from the time series of the state and that of the bifurcation parameter. Then the LSTM layer is used to identify long-term dependencies from the sequences of local features for tipping points prediction. We assume that the DL algorithm can detect features that emerge in time series prior to a tipping point, such as the features of the recovery rate in normal forms, which are associated with the occurrence of tipping points. We apply the DL algorithm to predict the occurrence of tipping points in systems it was not trained on. Our DL algorithm is applicable to both regularly-sampled and irregularly-sampled time series. We first test this DL algorithm on regularly-sampled time series generated by the models from ecology(28, 29, 30) and climatology(31, 32, 33, 34). The DL algorithm outperforms traditional methods in prediction accuracy for regularly-sampled model time series. We further validate our DL algorithm on irregularly-sampled time series generated by the same models and two other models from neuroscience with hysteresis phenomena(35, 36). Finally, we validate our DL algorithm in irregularly-sampled empirical data from microbiology(37), thermoacoustics(20) and paleoclimatology(38). In this work, we show that our DL algorithm is effective in dealing with dynamical systems exhibiting fold, Hopf and transcritical bifurcation. We anticipate that our DL algorithm may also be applicable to dynamical systems exhibiting other bifurcation types.
@@ Results
Fig.1   (A-C) The normal forms of fold, (supercritical) Hopf, and transcritical bifurcation where bifurcation occurs at $\mu=0$ (hollow square). Their mathematical expressions are inside the boxes within the corresponding figures. (D-F) The recovery rate $\lambda$ of the normal forms for the fold, (supercritical) Hopf, and transcritical bifurcation as a function of the control parameter $\mu$, their function expressions are inside the boxes within the corresponding figures. Bifurcation occurs when the recovery rate $\lambda$ reaches zero (hollow square).
@@
We applied the DL model on regularly-sampled time series generated from three ecological models with white noise and three climate models with red noise. We generated test data with 11 different initial values of the bifurcation parameter for each model and 50 test time series for each initial value of the bifurcation parameter. To evaluate the performance of DL algorithm, we measured the relative error(25) of tipping points prediction for each of the 50 test time series associated with every initial value of the bifurcation parameter. The performance of DL algorithm is evaluated by the mean relative error of tipping points prediction, which is averaged over the 50 measurements of the relative error of tipping points prediction for each initial value of the bifurcation parameter. We first compared the results of DL algorithm with those of degenerate fingerprinting(9), DEV(12) and null model in three ecological models with white noise which exhibit fold, Hopf and transcritical bifurcations respectively. Then we compared the results of DL algorithm with those of BB method(10), DEV(12) and null model in three climate models with red noise which exhibit fold, Hopf and transcritical bifurcations respectively. As Fig. 2 shows, the DL algorithm outperforms the other competing algorithms for all initial values of each model. Moreover, the DL algorithm exhibits smaller fluctuations of relative error of tipping points prediction than the other competing algorithms for each initial value and exhibits smaller fluctuations of the mean relative error of tipping points prediction than the other competing algorithms across different initial values. Our results suggest that the DL algorithm is more robust against different initial values of the bifurcation parameter than competing algorithms.
Fig.2   We tested the DL model on regularly-sampled model time series with 11 initial values of bifurcation parameter. Here the mean relative error of tipping points prediction is plotted as a function of initial parameter. The area covered by the polyline represents the 90% confidence interval for the relative error of tipping points prediction. (a-c) We compared the DL algorithm (red lines) with degenerate fingerprinting (blue lines), DEV (green lines) and null model (purple lines) on three ecological model time series with white noise. These model time series undergo fold, Hopf, and transcritical bifurcation, occurring at $0.268$, $0.480$, and $5.882$, respectively. (d-f) The DL algorithm (red lines) is compared with BB method (blue lines), DEV (green lines) and null model (purple lines) on three climate model time series with red noise. These model time series undergo fold, Hopf, and transcritical bifurcation, occurring at $0.962$, $0.355$, and $-0.005$, respectively.
@@
We further tested the DL model on irregularly-sampled time series from the above six models. It is worth noting that the above competing algorithms are not applicable for irregularly-sampled time series. Therefore, we used linear interpolation to transform these irregularly-sampled time series into equidistant data. This allows us to use the competing algorithms of degenerate fingerprinting, BB method and DEV for detecting early warning signals on reconstructed model time series(42). The null model is also applied as a competing algorithm. From Fig. 3, we find that DL algorithm outperforms the other competing algorithms for all initial values of each model. Moreover, our results suggest that the DL algorithm is more robust against different initial values of the bifurcation parameter than competing algorithms.
@@
Fig.3   We tested the DL model on irregularly-sampled model time series with 11 initial values of bifurcation parameter. Here the mean relative error of tipping points prediction is plotted as a function of initial parameter. The area covered by the polyline represents 90% confidence interval for the relative error of tipping points prediction. (a-c) We compared the DL algorithm (red lines) with degenerate fingerprinting (blue lines), DEV (green lines) and null model (purple lines) on three ecological model time series with white noise. (d-f) The DL algorithm (red lines) is compared with BB method (blue lines), DEV (green lines) and null model (purple lines) on three climate model time series with red noise. We used linear interpolation to transform these irregularly-sampled time series into equidistant data so that they are suitable for degenerate fingerprinting, BB method and DEV.
@@
We then tested the DL model on irregularly-sampled time series from the ascending arousal system with a fold/fold hysteresis loop. We conducted an irregular sampling of 400 points within the bifurcation parameter intervals of $[0.1,0.7]$ and $[1.3,1.9]$ (bifurcation parameter decreases in this interval) respectively. We applied the DL model on these irregularly-sampled time series and predicted that the tipping points occur at 1.157 and 0.877 with relative errors of tipping points prediction of 0.84% and 1.49% respectively, as shown in Fig. 4 (a). Similarly, we tested the DL model on irregularly-sampled time series from the Sprott B system with a Hopf/Hopf-hysteresis bursting. We irregularly sampled 400 points from the bifurcation parameter intervals of $[\pi,1.3\pi]$ and $[1.7\pi,2\pi]$ (bifurcation parameter decreases in this interval) respectively. We applied the DL model on these irregularly-sampled time series and predicted that the tipping points occur at $1.461\pi$ and $1.540\pi$ with relative errors of tipping points prediction of 0.05% and 0.81% respectively, as shown in Fig. 4 (b). Our results suggest that the DL algorithm is effective for tipping point prediction in theoretical models with hysteresis phenomenon.
@@
Fig.4   The performance of the DL algorithm in predicting transitions of irregularly-sampled time series generated by theoretical models exhibiting hysteresis. The red and blue curves illustrate the time series forced by increasing and decreasing bifurcation parameter respectively, while the grey area denotes the range of time series data used for prediction. (a) In the case of the sleep-wake system, a fold/fold hysteresis loop is observed at $D=1.152$ and $D=0.883$. The predicted tipping points are $1.157$ (red dashed line) and $0.877$ (blue dashed line) respectively. (b) For the Sprott B system, a Hopf/Hopf hysteresis bursting occurs at $k=1.461\pi$ and $k=1.539\pi$. The predicted tipping points are $1.461\pi$ (red dashed line) and $1.540\pi$ (blue dashed line) respectively.
@@
We tested the DL model on three empirical examples, including a cyanobacteria microcosm experiment under light stress(37), a physical experiment of voice oscillation during heat(20) and metallic elements concentration in sediment cores collected from the eastern Mediterranean Sea(38). For each empirical system, we applied the DL model to four irregularly-sampled time series, each containing 400 data points, across different parameter ranges. Then we used linear interpolation to transform these irregularly-sampled records to time series with equidistant data. This allows us to use the competing algorithms of degenerate fingerprinting, BB method and DEV for detecting early warning signals on reconstructed empirical records(42). Note that the first two empirical examples from ecology and physics are subject to white noise while the third empirical example from paleoclimatology is subject to red noise(43). Therefore, we applied degenerate fingerprinting to the first two empirical examples and BB method to the third empirical example. We compared the results of the DL algorithm with those of four competing algorithms, including degenerate fingerprinting, BB method, DEV and null model.
[…]
@@
The second empirical data we analyzed is the thermoacoustic system. The state of the thermoacoustic system is measured by acoustic pressure. The thermoacoustic system undergoes a Hopf bifurcation from a non-oscillatory to an oscillatory state with increasing voltage(24, 44, 45, 46). It has also been demonstrated that the transition to high amplitude limit cycle oscillations occurs later for faster changing rate of voltage, which can be explained by rate-delayed tipping(20, 21). In order to verify the robustness of DL algorithm in tipping points prediction against different changing rates of voltage, we applied the DL model to empirical time series under three voltage changing rates (20mV/s, 40mV/s, 60mV/s) within the bifurcation parameter voltage ranges from 0 to 1 V, from 0.05 to 1.05 V, from 0.1 to 1.1 V and from 0.15 to 1.15V respectively. From Fig. 5 B1-B4, we find that the DL algorithm outperforms the other competing algorithms in predicting the tipping points, where the acoustic pressure starts to oscillate, for various voltage intervals of empirical data under the voltage changing rate of 20mV/s. We further test these algorithms for different changing rates of voltage as shown in Supplementary Fig. 16, which suggest the DL algorithms also shows best performance among all these algorithms. These results suggest that the DL algorithm is robust against various voltage intervals and changing rages of voltage.
The third empirical data is the concentration of metallic elements in sediment cores collected from the eastern Mediterranean Sea. The concentration of metallic elements changes rapidly due to the transition between oxic and anoxic states occurred regularly in this region. The bifurcation types of these transitions have been uncovered by previous study(18). We studied the Mo concentration in sediment cores and the U concentration in sediment cores collected in different periods of time, which are from 160 to 125 ka BP and from 300 to 240 ka BP respectively. We lacked information on the bifurcation parameter series data in this example, therefore we used age as a substitute for the bifurcation parameter. Since the Mo concentration in sediment cores undergoes fold bifurcation, we applied the DL model on this data within four age intervals, from 160 to 140 ka BP, from 159 to 139 ka BP, from 158 to 138 ka BP and from 157 to 137 ka BP respectively. However, the U concentration in sediment cores undergoes transcritical bifurcation. We applied the DL model on this data within four age intervals, from 300 to 268 ka BP, from 298 to 266 ka BP, from 296 to 264 ka BP and from 294 to 262 ka BP respectively. We find that the DL algorithm outperforms the other competing algorithms in predicting tipping points where the concentration of metallic elements Mo or U in sediment cores change dramatically for various age intervals of empirical data as shown in Fig. 5 C1-C4 and Fig. 5 D1-D4. These results also suggest the robustness of the DL algorithm in tipping points prediction against various age intervals.
Fig.5   The performance of DL algorithm in predicting tipping points on irregularly-sampled empirical time series. Grey region represents the range of empirical time series used for prediction. (A1-A4) Cyanobacterial population undergoes a fold bifurcation, which is measured by light attenuation coefficient. (B1-B4) The thermoacoustic system undergoes a Hopf bifurcation under the 20mV/s changing rate of the control parameter voltage. (C1-C4) The eastern Mediterranean Sea experienced an anoxic event, it exhibited a fold bifurcation, which was measured using the concentration of Mo in sediment cores. (D1-D4) The eastern Mediterranean Sea experienced another anoxia event, characterized by a transcritical bifurcation, which was measured using the concentration of U in sediment cores. We compared the performance of the DL algorithm (red dashed lines) with degenerate fingerprinting (blue dash-dot lines), BB method (cyan dash-dot lines), DEV (green dash-dot lines) and null model (purple dotted lines). We used linear interpolation to transform these irregularly-sampled time series into equidistant data so that they are suitable for degenerate fingerprinting, BB method and DEV.
Fig.6   The 2D CNN-LSTM architecture. The 2D convolutional kernel processes state series and bifurcation parameter series into one-dimensional local feature series. These features are then subjected to local max pooling, followed by analysis using an LSTM layer. Finally, the features processed by the LSTM are mapped to the prediction of tipping point, through a fully connected layer.

### v2

@@ Introduction
A classical mathematical tool to understand tipping points is bifurcation theory which focuses on how dynamical systems undergo sudden qualitative changes as a parameter crosses a threshold(7). There are basically two classes of methods based on bifurcation theory to deal with tipping points prediction. The first class of methods is using lag-1 autocorrelation which is based on “critical slowing down”, a paramount clue of whether a tipping point is approached(5, 8). Such methods are widely used in various complex systems, including degenerate fingerprinting(9), BB method(10) and ROSA(11). The second class of methods is using dynamical eigenvalue (DEV)(12) which is based on Takens embedding theorem(13). Yet, the success of the existing methods based on bifurcation theory are impaired by two fundamental limitations. First, these methods are not applicable for irregularly-sampled time series data which are commonly observed from various scientific fields, such as geoscientific measurements(14, 15), medical observations(16) and biological systems(17). Second, the performance of existing methods based on bifurcation theory in prediction accuracy is affected by several approximations used in these methods. One such approximation is that only the first-order term of the dynamical systems is considered in these methods while the impact of the higher-order terms is ignored. However, the higher-order terms become significant for prediction when a tipping point is approached(18) (Supplementary information Note S1). Besides, in the approximation of fast-slow systems(19), the delay on the bifurcation-tipping due to the changing rate of the bifurcation parameter is ignored(20, 21) (Supplementary information Note S2).
@@
Recent studies show that some machine learning based techniques have been used as effective early warning signals of tipping points by learning generic features of bifurcation(18, 22, 23, 24). However, they can not be used for predicting where the tipping points occur. A machine learning framework based on reservoir computing has been developed for tipping points prediction(25, 26, 27). This algorithm requires time series sampled from all interacting variables of study system for training and prediction. However, this information is often not available for real-world systems. Therefore, it is often not feasible to use this machine learning framework for predicting tipping points of real-world systems.
@@
To circumvent above limitations of existing methods for tipping points prediction, we develop a deep learning (DL) algorithm based on 2D CNN-LSTM architecture. Based on the embedding theorem for irregular sampling(28) (Supplementary information Note S3), this DL algorithm only requires the time series of the state and that of the bifurcation parameter from a single variable of a system (Fig. 1). The 2D CNN layer uses convolution kernels with length $d$ to extract features of the system reconstructed by a $d$-dimensional delay embedding. These convolutional kernels move along the time series, extracting features from each segment of the state time series and the bifurcation parameter time series, thereby generating sequences of feature representations. Then the LSTM layer is used to identify long-term dependencies from the sequences of features of the reconstructed system for tipping points prediction. We assume that the DL algorithm can detect features that emerge in time series prior to a tipping point, such as the features of the recovery rate in normal forms, which are associated with the occurrence of tipping points. We apply the DL algorithm to predict the occurrence of tipping points in systems it was not trained on. Our DL algorithm is applicable to both regularly-sampled and irregularly-sampled time series. We first test this DL algorithm on regularly-sampled time series generated by the models from ecology(29, 30, 31) and climatology(32, 33, 34, 35). The DL algorithm outperforms traditional methods in prediction accuracy for regularly-sampled model time series. We further validate our DL algorithm on irregularly-sampled time series generated by the same models and two other models from neuroscience with hysteresis phenomena(36, 37). Finally, we validate our DL algorithm in irregularly-sampled empirical data from microbiology(38) and thermoacoustics(20). In this work, we show that our DL algorithm is effective in dealing with dynamical systems exhibiting fold, Hopf and transcritical bifurcation. We anticipate that our DL algorithm may also be applicable to dynamical systems exhibiting other bifurcation types.
Fig. 1. The 2D CNN-LSTM architecture. The 2D convolutional kernel processes state series and bifurcation parameter series into one-dimensional features series of the reconstructed system $(x_{i},x_{i+1},\dots,x_{i+d-1})$. This features series is then subjected to local max pooling, followed by analysis using an LSTM layer. Finally, the features processed by the LSTM are mapped to the predicted tipping point.
@@ Results
Fig. 2. (A-C) The normal forms of fold, (supercritical) Hopf, and transcritical bifurcation where bifurcation occurs at $\mu=0$ (hollow square). Their mathematical expressions are inside the boxes within the corresponding figures. (D-F) The recovery rate $\lambda$ of the normal forms for the fold, (supercritical) Hopf, and transcritical bifurcation as a function of the control parameter $\mu$, their function expressions are inside the boxes within the corresponding figures. Bifurcation occurs when the recovery rate $\lambda$ reaches zero (hollow square).
@@
We applied the DL model on regularly-sampled time series generated from three ecological models with white noise and three climate models with red noise. We generated test data for each model by changing the bifurcation parameter with eleven different initial values and five different changing rates. We generated 10 test time series for each initial value and changing rate of the bifurcation parameter (50 test time series for each initial value). To evaluate the performance of DL algorithm, we measured the relative error(25) of tipping points prediction for each of the 50 test time series associated with every initial value of the bifurcation parameter. The performance of DL algorithm is evaluated by the mean relative error of tipping points prediction, which is averaged over the 50 measurements of the relative error of tipping points prediction for each initial value of the bifurcation parameter. We designed an ablation study by dropping out the 2D CNN layer, with LSTM serving as a competing algorithm. We first compared the results of DL algorithm with those of degenerate fingerprinting(9), DEV(12) and LSTM in three ecological models with white noise which exhibit fold, Hopf and transcritical bifurcations respectively. Then we compared the results of DL algorithm with those of BB method(10), DEV(12) and LSTM in three climate models with red noise which exhibit fold, Hopf and transcritical bifurcations respectively. As Fig. 3 shows, the DL algorithm outperforms the other competing algorithms for all initial values of each model. Moreover, the DL algorithm exhibits smaller fluctuations of relative error of tipping points prediction than the other competing algorithms for each initial value and exhibits smaller fluctuations of the mean relative error of tipping points prediction than the other competing algorithms across different initial values. Our results suggest that the DL algorithm is more robust against different initial values of the bifurcation parameter than competing algorithms.
@@
We further tested the DL model on irregularly-sampled time series from the above six models. It is worth noting that the above competing algorithms are not applicable for irregularly-sampled time series. Therefore, we used linear interpolation to transform these irregularly-sampled time series into equidistant data. This allows us to use the competing algorithms of degenerate fingerprinting, BB method and DEV for detecting early warning signals on reconstructed model time series(42). LSTM is also applied as a competing algorithm. From Fig. 4, we find that DL algorithm outperforms the other competing algorithms for all initial values of each model. Moreover, our results suggest that the DL algorithm is more robust against different initial values of the bifurcation parameter than competing algorithms.
@@
Fig. 3. We tested the DL model on regularly-sampled model time series with eleven initial values of bifurcation parameter. Here the mean relative error of tipping points prediction is plotted as line graphs against the initial value of the parameter. The area covered by the polyline represents the 90% confidence interval for the relative error of tipping points prediction. (a-c) We compared the DL algorithm (red lines) with degenerate fingerprinting (blue lines), DEV (green lines) and LSTM (purple lines) on three ecological model time series with white noise. These model time series undergo fold, Hopf, and transcritical bifurcation, respectively. (d-f) The DL algorithm (red lines) is compared with BB method (blue lines), DEV (green lines) and LSTM (purple lines) on three climate model time series with red noise. These model time series undergo fold, Hopf, and transcritical bifurcation, respectively.
In the ablation study, our DL algorithm achieves a mean relative error about 9% in predicting tipping points on these model time series, while LSTM exhibits a mean relative error about 23%, as shown in Supplementary information Fig. S1 and Fig. S2. These results suggest that the accuracy of predicting tipping points is higher when the LSTM layer receives sequences of features from the reconstructed system, rather than sequences of a single variable of the system. We also plotted the line graphs of the mean relative error of tipping points prediction against the distance between the final value of bifurcation parameter time series and the value of the tipping point, as shown in Supplementary information Fig. S3 and Fig. S4. Our results suggest that the DL algorithm is more robust against different distance to the tipping point than competing algorithms. Moreover, we plotted the line graphs of the mean relative error of tipping points prediction against the initial value of the parameter in five different changing rates of the bifurcation parameter, as shown in Supplementary information Fig. S5 and Fig. S6. Our results suggest that the DL algorithm is robust against various changing rates of bifurcation parameter.
Fig. 4. We tested the DL model on irregularly-sampled model time series with eleven initial values of bifurcation parameter. Here the mean relative error of tipping points prediction is plotted as line graphs against the initial value of the parameter. The area covered by the polyline represents 90% confidence interval for the relative error of tipping points prediction. (a-c) We compared the DL algorithm (red lines) with degenerate fingerprinting (blue lines), DEV (green lines) and LSTM (purple lines) on three ecological model time series with white noise. (d-f) The DL algorithm (red lines) is compared with BB method (blue lines), DEV (green lines) and LSTM (purple lines) on three climate model time series with red noise. We used linear interpolation to transform these irregularly-sampled time series into equidistant data so that they are suitable for degenerate fingerprinting, BB method and DEV.
@@
We then tested the DL model on irregularly-sampled time series from the ascending arousal system with a fold/fold hysteresis loop. We conducted irregular sampling of 400 points from twenty-two initial values of the bifurcation parameter, which are $0.1,0.2,\dots,1.1$ (bifurcation parameter increases from these initial values) and $1.9,1.8,\dots,0.9$ (bifurcation parameter decreases from these initial values). We applied the DL model on these irregularly-sampled time series and the mean relative errors of tipping points prediction are 3.12% and 3.31% respectively, for the increasing and decreasing bifurcation parameter cases, as shown in Fig. 5 (a). Similarly, we tested the DL model on irregularly-sampled time series from the Sprott B system with a Hopf/Hopf-hysteresis bursting. We irregularly sampled 400 points from twenty-two initial values of the bifurcation parameter, which are $\pi,1.04\pi,\dots,1.4\pi$ (bifurcation parameter increases from these initial values) and $2\pi,1.96\pi,\dots,1.6\pi$ (bifurcation parameter decreases from these initial values). We applied the DL model on these irregularly-sampled time series and the mean relative errors of tipping points prediction are 2.82% and 2.54% respectively, for the increasing and decreasing bifurcation parameter cases, as shown in Fig. 5 (b). Our results suggest that the DL algorithm is effective for tipping point prediction in theoretical models with hysteresis phenomenon.
@@
Fig. 5. The performance of the DL algorithm in predicting transitions of irregularly-sampled time series generated by theoretical models exhibiting hysteresis. The red and blue curves illustrate the time series forced by increasing and decreasing bifurcation parameter respectively. The orange dots denote the DL predictions, and the orange short vertical lines denote the initial points of the time series data used for prediction. We connect them with orange lines. (a) In the case of the sleep-wake system, a fold/fold hysteresis loop is observed at $D=1.153$ (red dashed line) and $D=0.883$ (blue dashed line) respectively. The predicted tipping points are between $[1.148,1.169]$ and $[0.8679,0.8898]$ respectively. (b) For the Sprott B system, a Hopf/Hopf hysteresis bursting occurs at $k=1.461\pi$ (red dashed line) and $k=1.539\pi$ (blue dashed line) respectively. The predicted tipping points are between $[1.454\pi,1.464\pi]$ and $[1.534\pi,1.545\pi]$ respectively.
@@
We tested the DL model on two empirical examples, including a cyanobacteria microcosm experiment under light stress(38) and a physical experiment of voice oscillation during heat(20). For each empirical system, we applied the DL model to several irregularly-sampled time series. Each time series includes 400 data points with different initial values of the control parameter. Then we used linear interpolation to transform these irregularly-sampled records to time series with equidistant data. This allows us to use the competing algorithms of degenerate fingerprinting, BB method and DEV for detecting early warning signals on reconstructed empirical records(42). Note that these two empirical examples from ecology and physics are subject to white noise. Therefore, we applied degenerate fingerprinting to these empirical data and did not use the BB method for comparison. We compared the results of the DL algorithm with two competing algorithms, which are degenerate fingerprinting and DEV.
In the cyanobacteria microcosm experiment under light stress, the photo-inhibition drives a cyanobacterial population, measured by light attenuation coefficient, towards a tipping point with fold bifurcation when a critical threshold of light irradiance is approached(38). We sampled seven time series with different initial values of light irradiance which are 477, 517, 557, 903, 944, 985 and 1025 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$. Since the data between 879 and 903 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$ is missing, the time series is sampled with initial values of 477, 517, 557 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$ respectively and the same final value of 879 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$. The mean relative error of the prediction with our DL algorithm is 3.82%, where the ground truth of the tipping point is 1091 $\rm\mu$mol photons $\rm m^{-2}s^{-1}$. As shown in Fig. 6 A1-A3, the DL algorithm outperforms the other competing algorithms in all cases. These results suggest the robustness of the DL algorithm in tipping points prediction against different initial values of light irradiance.
@@
The second empirical data we analyzed is the thermoacoustic system. The state of the thermoacoustic system is measured by acoustic pressure. The thermoacoustic system undergoes a Hopf bifurcation from a non-oscillatory to an oscillatory state with increasing voltage(24, 43, 44, 45). It has also been demonstrated that the transition to high amplitude limit cycle oscillations occurs later for faster changing rate of voltage, which can be explained by rate-delayed tipping(20, 21). We sampled time series with nine initial values of voltage which are $0,0.2,\dots,1.6$ V under voltage changing rates of 20 mV/s and 40 mV/s, and with ten initial values of voltage which are $0,0.2,\dots,1.8$ V under a voltage changing rate of 60 mV/s. We then applied a change-point detection algorithm(46) to detect tipping points of these empirical time series under voltage changing rates of 20 mV/s, 40 mV/s and 60 mV/s, which are 1.72 V, 1.76 V and 1.87 V respectively. As shown in Fig. 6 B1-B3, we find that the DL algorithm outperforms the other competing algorithms in predicting the tipping points in most cases under the voltage changing rate of 40 mV/s, with mean relative error of 4.31% by our DL algorithm. We further test these algorithms for other two changing rates of voltage as shown in Supplementary information Fig. S7, which suggests the DL algorithms also shows best performance among all these algorithms. These results suggest that the DL algorithm is robust against various initial values and changing rages of voltage.